### v1

# Multiplier ideals of normal surface singularities

Abstract.
We study the multiplier ideals and the corresponding jumping numbers and multiplicities $\{m(c)\}_{c\in\mathbb{R}}$ in the following context: $(X,o)$ is a complex analytic normal surface singularity, ${\mathfrak{a}}\subset{\mathscr{O}}_{X,o}$ is an ${\mathfrak{m}}_{X,o}$–primary ideal, $\phi:\widetilde{X}\to X$ is a log resolution of $\mathfrak{a}$ such that $\mathfrak{a}{\mathscr{O}}_{\widetilde{X}}={\mathscr{O}}_{\widetilde{X}}(-F)$, for some nonzero effective divisor $F$ supported on $\phi^{-1}(0)$.
We show that $\{m(c)\}_{c>0}$ is combinatorially computable from $F$ and the resolution graph $\Gamma$ of $\phi$, and we provide several formulae. We also extend Budur’s result (valid for $(X,o)=(\mathbb{C}^{2},0)$), which makes an identification of $\sum_{c\in[0,1]}m(c)t^{c}$ with a certain Hodge spectrum. In our general case we use Hodge spectrum with coefficients in a mixed Hodge module.
We show that $\{m(c)\}_{c\leq 0}$ usually depends on the analytic type of $(X,o)$. However, for some distinguished analytic types we determine it concretely. E.g., when $(X,o)$ is weighted homogeneous (and $F$ is associated with the central vertex), we recover $\sum_{c}m(c)t^{c}$ from the Poincaré series of $(X,o)$ and when $(X,o)$ is a splice quotient then we recover $\sum_{c}m(c)t^{c}$ from the multivariable topological Poincaré (zeta) function of $\Gamma$.
Key words and phrases:
normal surface singularities, plane curve singularities, multiplier ideals, jumping numbers, jumping multiplicities, Hodge spectrum, mixed Hodge modules, canonical covering, weighted homogeneous singularities, splice quotient singularities
2010 Mathematics Subject Classification:
Primary: 14B05, 14Fxx, 32S05; Secondary: 32S10, 32S25
The authors are partially supported by NKFIH Grant “Élvonal (Frontier)” KKP 144148. The second author acknowledges the support of the ’János Bolyai Research Scholarship’ of the Hungarian Academy of Sciences, and the support of the project “Singularities and Applications” - CF 132/31.07.2023 funded by the European Union - NextGenerationEU - through Romania’s National Recovery and Resilience Plan.
1. Introduction
Assume that $(X,o)$ is a complex analytic normal surface singularity, let ${\mathscr{O}}_{X,o}$ be its local algebra with maximal ideal ${\mathfrak{m}}_{X,o}$. Let us also fix an ${\mathfrak{m}}_{X,o}$–primary ideal ${\mathfrak{a}}\subset{\mathscr{O}}_{X,o}$. Then there exists a log resolution $\phi:\widetilde{X}\to X$ and a nonzero integral cycle $F$ supported on the exceptional curve $\phi^{-1}(0)=E$ such that ${\mathfrak{a}}{\mathscr{O}}_{\widetilde{X}}={\mathscr{O}}_{\widetilde{X}}(-F)$. In fact, in our study we will start with some $\phi$ and such a cycle $F$.
We denote by $Z_{K}$ the anticanonical cycle associated with $\phi$, $-c_{1}(\Omega^{2}_{\widetilde{X}})=-c_{1}({\mathscr{O}}_{\widetilde{X}}(K_{\widetilde{X}}))\in H^{2}(\widetilde{X},\mathbb{Z})$, identified with a rational cycle (via the adjunction relations). We consider for any $c\in\mathbb{R}$ the cycle $L(c):=\lfloor Z_{K}+cF\rfloor$. Then the tower of multiplier ideals is defined as $\mathfrak{I}(c,F):=\phi_{*}{\mathscr{O}}_{\widetilde{X}}(-L(c))\subset{\mathscr{O}}_{X,o}$. They measure the complexity of the singularity $(X,o)$ compatibly with $F$ (or with $\mathfrak{a}$). They were introduced in [Ein, Laz2] and since then they were intensively studied, see [budur, Carrami, Carrami2, Galindo, Galindo2, Howald, Hyry, Jarvilehto, Montaner, Naie, Tucker] with all their connections with different aspects of singularities and singular map-germs. In fact, the definition and study extends to any dimension, see e.g. [Ein, Laz2, Pande] and the references therein.
It turns out that for any $c\in\mathbb{R}$, the ideal $\mathfrak{I}(c,F)\subset{\mathscr{O}}_{X,o}$ has finite codimension, $\mathfrak{I}(c,F)\subset\mathfrak{I}(c^{\prime},F)$ for $c^{\prime}<c$, and $\mathfrak{I}(c,F)={\mathscr{O}}_{X,o}$ for $c\ll 0$. The number $c$ is called jumping number if $m(c):=\dim\,\mathfrak{I}(c-\epsilon,F)/\mathfrak{I}(c,F)\not=0$ for $0<\epsilon\ll 1$, and $m(c)$ is the corresponding jumping multiplicity. One verifies that in any compact interval there are only finitely many jumping numbers and they are all rational.
Compared with the earlier literature in the subject we make two major extensions:
(A)  We do not assume any restriction regarding the surface singularity $(X,o)$, e.g. we do not assume that the link is a rational homology sphere (i.e. the dual graph might have loops and the irreducible exceptional divisors might have higher genera). Also, we do not restrict the analytic type of $(X,o)$: on the supporting topological type the analytic type might vary (which in the non-rational or non-minimally elliptic cases might have considerable effects). We emphasize that in general the classification of such analytic types (or the description of the corresponding moduli spaces) is still an open problem. Under this generality the general analytic tools (e.g. vanishing theorems) are more complex (and the zones what they do not cover are wider).
(B)  In the earlier studies only the cases $c>0$ were treated. Our study of the cases $c\leq 0$ opens a new perspectives. If $(X,o)$ is smooth or log terminal then $m(c)=0$ for $c\leq 0$, however, in the general case (even for rational $(X,o)$) the negative zone appears as a very challenging part.
In fact (and this is one of the main messages of the note), we show that the study of the Poincaré series of multiplier ideals $\sum_{c\in\mathbb{R}}m(c)t^{c}$ (introduced first in [Galindo]) breaks into two major parts with rather different aspects.
(i)  (the $c>0$ zone)  We also call this the Riemann–Roch zone, since in this case we determine $\sum_{c>0}m(c)t^{c}$ completely topologically from the resolution graph $\Gamma$ of $\phi$ and the cycle $F$ (without any restriction regarding the link or the analytic type of $(X,o)$).
We provide several combinatorial formulae. The first one, Proposition LABEL:lem:chi, is a generalization of expressions obtained in [Carrami2, Carrami, Tucker] proved in those articles for rational singularities.
We also compute $m(c+1)-m(c)$ (see Corollary LABEL:cor:m0), thus we recover $\sum_{c>0}m(c)t^{c}$ from $\sum_{c\in(0,1]}m(c)t^{c}$ and $(D_{c},F)$, where $D_{c}=L(c)-L(c-\epsilon)$ ($0<\epsilon\ll 1$) is the jumping divisor (introduced in [Carrami2]), computable combinatorially. (This fact was also shown for rational $(X,o)$ in [Carrami].)
Hence, the first interval and $\sum_{c\in(0,1]}m(c)t^{c}$ has a distinguished role.
Regarding this we stress out two facts. Firstly, Budur in [budur] identified $\sum_{c\in(0,1]}m(c)t^{c}$ with the Hodge spectrum ${\rm Sp}_{(0,1]}(f)$ of a function germ $f$, where in his case, $(X,o)=(\mathbb{C}^{2},0)$ is smooth, $f:(\mathbb{C}^{2},0)\to(\mathbb{C},0)$ is a function germ with an isolated singularity, $\phi$ is a log embedded resolution and $F$ is the divisor ${\rm div}_{\phi,E}(f)$ of $\phi^{*}(f)$ along $\phi^{-1}(0)=E$. This was reproved in [Carrami] as well. (For the theory of mixed Hodge structures and of the spectrum associated with isolated hypersurface singularities see the work of Steenbrink, e.g. [steenbrink, steenbrink2] or [SSS].)
In this note we succeed to generalize this statement. The first step is that we rewrite $J_{[0,1]}(t)=\sum_{c\in[0,1]}m(c)t^{c}$ in a different way, in terms of $\Gamma$ and the multiplicities of $F$, for any $(X,o)$ and any $F$ (cf. Theorem LABEL:th:main_mult). This is the second combinatorial formula of the jumping multiplicities.
This formula will resonate perfectly with Hodge spectrum computations of a generic element $f\in\mathfrak{a}$. Indeed, if $(X,o)$ is numerically Gorenstein (that is, $Z_{K}$ is an integral cycle), and $f$ is an analytic function gems $(X,o)\to(\mathbb{C},0)$ with isolated singularities, ${\rm div}_{\phi,E}(f)=F$, we prove the generalization of the statement of Budur: $J_{[0,1]}(t)={\rm Sp}_{[0,1]}(t)$, cf. Corollary LABEL:cor:NS. However, when $\Gamma$ is not numerically Gorenstein, $\{Z_{K}\}\not=0$, then we need to use a more general spectrum computation established in [NS]. In [NS] the usual spectrum is replaced by the spectrum of $f$ with coefficients in a mixed Hodge module ${\mathscr{M}}$ of $(X,o)$, such that ${\mathscr{M}}$ restricted to $X\setminus\{o\}$ is a variation of Hodge structure whose underlying representation is abelian. Indeed, under the assumption that the link is a rational homology sphere, from $Z_{K}$ we construct a character of the first homology of the link, hence a rank one variation of Hodge structure ${\mathbb{V}}$ on $X\setminus\{o\}$, and we prove the identity $J_{[0,1]}(t)={\rm Sp}_{[0,1]}(t)(f,j_{*}{\mathbb{V}})$ (where $j:X\setminus\{o\}\subset X$ is the inclusion), see Theorem LABEL:th:NS2. The construction of ${\mathbb{V}}$ can be done using the canonical cyclic covering of $(X,o)$ (associated with $Z_{K}$, or in $\mathbb{Q}$–Gorenstein case associated with the canonical divisor $K_{\widetilde{X}}$, as well). We emphasize again, though a priori both objects $J_{[0,1]}(t)$ and ${\rm Sp}_{[0,1]}(t)(f,j_{*}{\mathbb{V}})$ are analytical, we provide combinatorial formulae for them.
The second new aspect regarding $\sum_{c\in(0,1]}m(c)t^{c}$ is the following connection. Assume that there exists again a function $f:(X,o)\to(\mathbb{C},0)$ with an isolated singularity such that $F={\rm div}_{\phi,E}(f)$. Set $(C,o):=(f^{-1}(0),o)$ and let $\delta(C,o)$ be its delta invariant. Then $\sum_{c\in(0,1]}m(c)=\delta(C,o)+(\{Z_{K}\},F)$.
(ii)  (the $c\leq 0$ zone)  We show by examples that the multiplicity $m(0)$ cannot be determined topologically. However, we characterize it completely in terms of the analytic type of $(X,o)$. The analytic contribution is a term $\mathfrak{h}^{0}\in\{0,1\}$ characterized by an $h^{0}$–sheaf cohomology.
The $c\leq 0$ zone can be called the ‘$h^{0}$–zone’, where the Riemann–Roch expressions or vanishing theorems cannot be applied. In this zone $m(c)$ usually cannot be computed form $\Gamma$ and $F$. We exemplify this fact by concrete examples when we fix $\Gamma$ and $F$ and we consider several different analytic realizations, and we compute the $m(c)$’s. Interestingly enough, already the numerical invariant $\sum_{c\leq 0}m(c)$ measures the analytic type and it is connected to the most important analytic invariant of $(X,o)$, the geometric genus $p_{g}(X,o)$. More precisely, we prove that $\sum_{c\leq 0}m(c)=p_{g}(X,o)+\chi(\lfloor Z_{K}\rfloor)$. Already when $(X,o)$ is rational (i.e. $p_{g}(X,o)=0$), $\sum_{c\leq 0}m(c)$ can be arbitrarily large (see e.g. Example LABEL:ex:sum).
On the other hand, even for this zone, we provide concrete formulae valid for certain distinguished analytic types.
$\bullet$  In the case of a weighted homogeneous germs $(X,o)$ with local graded algebra $\oplus_{\ell\geq 0}R_{\ell}$ (and $F$ is associated with the central vertex), we show the correspondence $\dim(R_{\ell})=m(c)$ with $c=(\ell-\mathfrak{r})/\alpha$ (for notations and precise statement see subsection LABEL:ss:we). For $\ell>{\mathfrak{r}}$ (corresponding to $c>0$) $\dim(R_{\ell})$ is topological computable from the Seifert invariants of the link, however in general it is not. But, if the link is a rational homology sphere, then by Dolgachev-Pinkham-Demazure formulae all the terms for any $\ell$ are topological. (For the theory of weighted homogeneous germs see e.g. [pinkham].)
$\bullet$  If the link is a rational homology sphere and we fix a resolution graph $\Gamma$, and we assume that the analytic type of $(X,o)$ is splice quotient with respect to $\Gamma$ (a family introduced by Neumann and Wahl, cf. [NW]). Then in Proposition LABEL:PZ we provide a combinatorial formula for $\sum_{c\in\mathbb{R}}m(c)t^{c}$ in terms of $\{L(c)\}_{c}$ and the multivariable topological Poincaré (zeta) function $Z({\bf t})$ of $\Gamma$. This includes e.g. all the cases when $(X,o)$ is rational or minimally elliptic (with $\Gamma$ minimal). (For more on $Z({\bf t})$ see [CDG, CDGEq, Nkonyv].)
But we emphasize again that in general, the computation of $\{m(c)\}_{c\leq 0}$ is of analytic nature, hence it runs in parallel with the classification of the analytic types of $(X,o)$ supported of a fixed topological type identified by $\Gamma$. We will return back to this study in the forthcoming manuscript.
2. Preliminaries
2.1. Notations regarding resolutions and resolution graphs
2.1.1.
Let us fix a normal surface singularity $(X,o)$ and one of its log (good) resolutions $\phi:\widetilde{X}\to X$. Let $\Gamma$ be the dual resolution graph of $\phi$, we denote its vertices by $\mathscr{V}$ and its edges by $\mathscr{E}$. We write $E$ for the irreducible exceptional curve $\phi^{-1}(o)$, and let $\{E_{i}\}_{i\in\mathscr{V}}$ be its irreducible components.
The lattice $L:=H_{2}(\widetilde{X},\mathbb{Z})$ is a free $\mathbb{Z}$–module generated by the classes of $\{E_{i}\}_{i\in\mathscr{V}}$ and it is endowed with the natural negative definite intersection form (,)(\,,\,)( , ). The dual lattice is $L^{\prime}={\rm Hom}_{\mathbb{Z}}(L,\mathbb{Z})\simeq\{l^{\prime}\in L\otimes\mathbb{Q}\,:\,(l^{\prime},L)\in\mathbb{Z}\}\subset L\otimes\mathbb{Q}$. We use the same symbol (,)(\,,\,)( , ) for the extended intersection form to $L\otimes\mathbb{Q}$. The overlattice $L^{\prime}$ is generated over $\mathbb{Z}$ by the (anti)dual classes $\{E^{*}_{i}\}_{i\in\mathscr{V}}$ defined by $(E^{*}_{i},E_{j})=-\delta_{ij}$ (where $\delta_{ij}$ stays for the Kronecker symbol). $L^{\prime}$ can also identified with $H^{2}(\widetilde{X},\mathbb{Z})$ (the target of the first Chern class $c_{1}:{\rm Pic}(\widetilde{X})\to H^{2}(\widetilde{X},\mathbb{Z})$).
We denote by $M$ the link of $(X,o)$. It is known that $M$ is a rational homology sphere if and only if each $E_{i}$ is rational and $\Gamma$ is a tree. It is an integral homology sphere if additionally $\det\,(\,,\,)=\pm 1$. More generally, if $g_{i}$ denotes the genus of $E_{i}$, $g:=\sum_{i}g_{i}$ and $h$ the first Betti number of the topological realization of $\Gamma$ (the number of independent cycle in the graph), then the first Betti number of $M$ is $2g+h$, and ${\rm Tors}\,H_{1}(M,\mathbb{Z})\simeq L^{\prime}/L$.
There is a natural partial ordering of $L\otimes\mathbb{Q}$: we write $l_{1}^{\prime}\geq l_{2}^{\prime}$ if $l_{1}^{\prime}-l_{2}^{\prime}=\sum_{i}a_{i}E_{i}$ with every $a_{i}\geq 0$. We set $L_{\geq 0}=\{l\in L\,:\,l\geq 0\}$ and $L_{>0}=L_{\geq 0}\setminus\{0\}$.
We define the rational Lipman cone as ${\mathscr{S}}_{\mathbb{Q}}:=\{l^{\prime}\in L\otimes\mathbb{Q}\,:\,(l^{\prime},E_{i})\leq 0\ \mbox{for all $i$}\}$, and we also set ${\mathscr{S}}^{\prime}:={\mathscr{S}}_{\mathbb{Q}}\cap L^{\prime}$ and ${\mathscr{S}}:={\mathscr{S}}_{\mathbb{Q}}\cap L$. As a monoid ${\mathscr{S}}^{\prime}$ is generated over ${\mathbb{Z}}_{\geq 0}$ by $\{E^{*}_{i}\}_{i}$. If $s^{\prime}\in{\mathscr{S}}_{\mathbb{Q}}\setminus\{0\}$, then all its $E_{i}$–coordinates are strictly positive (see [Nkonyv, Corollary 2.1.19]).
Let $K_{\widetilde{X}}$ be the canonical divisor of $\widetilde{X}$ (well–defined up to a linear equivalence), that is, ${\mathscr{O}}_{\widetilde{X}}(K_{\widetilde{X}})\simeq\Omega^{2}_{\widetilde{X}}$, the sheaf of holomorphic 2–forms. We denote its first Chern class $c_{1}({\mathscr{O}}_{\widetilde{X}}(K_{\widetilde{X}}))\in H^{2}(\widetilde{X},\mathbb{Z})\simeq L^{\prime}$ by $-Z_{K}$; $Z_{K}$ is called the (anti)canonical cycle of $\phi$. The (rational) cycle $Z_{K}\in L^{\prime}$ can also be determined combinatorially from (,)(\,,\,)( , ) by the adjunction formulae $(Z_{K},E_{i})=(E_{i},E_{i})+2-2g_{i}$ for all $i\in\mathscr{V}$.
We call $(X,o)$ numerically Gorenstein if $Z_{K}\in L$, i.e. if all its $E_{i}$–coefficients are integral (this condition is independent of the choice of $\phi$). We call $(X,o)$ Gorenstein if $Z_{K}\in L$ and one can choose $K_{\widetilde{X}}$ as $-Z_{K}$, i.e. ${\mathscr{O}}_{\widetilde{X}}(K_{\widetilde{X}}+Z_{K})\simeq{\mathscr{O}}_{\widetilde{X}}$. In a different language, $(X,o)$ is numerically Gorenstein (resp. Gorenstein) exactly when the line bundle $\Omega^{2}_{X\setminus\{o\}}$ is topologically (resp. analytically) trivial. (For a different characterization see also LABEL:bek:qg.)
Finally, we set for any $l^{\prime}\in L^{\prime}$ the Riemann-Roch expression $\chi(l^{\prime})=(Z_{K}-l^{\prime},l^{\prime})/2$. If $l\in L_{>0}$, then $\chi(l)=h^{0}({\mathscr{O}}_{l})-h^{1}({\mathscr{O}}_{l})$. E.g., $\chi(E)=1-g-h$. As usual, we also write $\chi({\mathscr{O}}_{l}(D))$ for $h^{0}({\mathscr{O}}_{l}(D))-h^{1}({\mathscr{O}}_{l}(D))$, which, by Riemann-Roch, equals $\chi(l)+(D,l)$ ($D\in L$).
For any ${\mathbb{R}}$-divisor $D=\sum_{i}d_{i}E_{i}$ of $\widetilde{X}$ ($d_{i}\in{\mathbb{R}}$) we define its integral part $\lfloor D\rfloor$ and fractional part $\{D\}$ by applying the corresponding operation to each coefficient $d_{i}$.
For more details see e.g. [NOSz, Nkonyv].